# Symmetry transformation group arising from the Laplace–Runge–Lenz vector

### v2

@@ 1. Introduction
The Kepler problem in Newtonian mechanics has been widely studied and encompasses both a particle moving under inverse-square force and two bodies moving under mutual gravitational attraction. One of its most interesting features is the existence of the conserved Laplace–Runge–Lenz (LRL) vector (see e.g. [1, 2] and references therein). It lies in the plane of motion, which is orthogonal to the angular momentum vector, and points in the direction given by the periapsis in the case of inverse-square motion and the position of closest separation in the case of two-body motion. This conserved vector has a long history, going back over three centuries, and plays a key role in understanding the shape and orientation of astronomical orbits in the solar system. Its quantum mechanical analog was important historically in the initial derivation of energy levels in the hydrogen atom (see e.g. [3] and references therein).
@@
In contrast to the conservation of energy and angular momentum in the Kepler problem, which come directly from Noether’s theorem using the manifest time-translation and rotation symmetries of Newtonian space, the LRL vector does not arise from any kinematic symmetry principle. Nevertheless, Noether’s theorem in reverse [4, 5] states that every conserved quantity is generated by some infinitesimal symmetry transformation that leaves invariant the action principle for the Kepler problem. Since the LRL vector comprises three independent conserved quantities, its symmetry generator consists of three independent infinitesimal transformations, which were first worked out in Ref.[6]. They turn out to involve the velocity variable, unlike the infinitesimal time-translation and rotation transformations which involve only the time and position variables. In particular, the LRL infinitesimal symmetry transformations are primary examples of what are called dynamical symmetries, in contradistinction to point symmetries which are typified by infinitesimal time-translations and rotations. The only additional point symmetry admitted by the Kepler problem consists of a scaling of the time and position variables, but this scaling transformation does not leave the action principle invariant.
@@
The five point symmetries together with the three LRL symmetries in the Kepler problem form an 8-dimensional Lie algebra in which the time-translation and scaling symmetries constitute a two-dimensional abelian subalgebra, namely they commute with the rotations and LRL symmetries. The latter symmetries are known to form a 6-dimensional Lie subalgebra when the LRL symmetries are normalized to have the same physical dimension as the rotations. This subalgebra’s structure turns out to be $\mathfrak{so}(4)$ when the energy is negative, and $\mathfrak{so}(3,1)$ when the energy is positive [7]. These Lie algebras are isomorphic to the symmetry algebra of isometries of, respectively, a 3-sphere $S^{3}$ in $\mathbb{R}^{4}$ and a 3-hyperboloid $H^{3}$ in Minkowski space $\mathbb{R}^{3+1}$.
@@
One explanation for the appearance of four-dimensional spaces is provided by Moser’s observation [8] that, in the negative energy case, the Kepler problem is equivalent to non-affine geodesic motion on $S^{3}$ in a four-dimensional auxiliary space under stereographic projection. A kinematical description of the projection is given by [9, 7] identifying the stereographic coordinates as the components of the Kepler velocity variable. The rotations in the auxiliary space correspond in the Kepler problem to changing the angular momentum while keeping the energy unchanged. These rotations constitute a $SO(4)$ transformation group that combines the ordinary physical $SO(3)$ rotations and the LRL symmetry transformations. A similar equivalence [3] holds in the positive energy case by considering non-affine geodesic motion on $H^{3}$.
@@
Another remaining question, which is not commonly discussed in the literature, arises from the magnitude of LRL vector. This quantity is given by a scalar expression in terms of the energy and magnitude of the angular momentum. The corresponding infinitesimal symmetry transformation, coming from Noether’s theorem in reverse, is just a linear combination of the generators of time-translation and rotation in the plane of motion. Namely it is a point symmetry, involving only the time and position variables. On the one hand, this choice of normalization of the LRL vector is essential for the subalgebra generated by the LRL symmetries and rotations to be $\mathfrak{so}(4)$ when energy is negative and $\mathfrak{so}(3,1)$ when energy is positive. Yet on the other hand, it is a free choice, which raises the question of whether a different or perhaps simpler structure is possible for the resulting symmetry subalgebra if some other choice of normalization is made. A natural choice is to normalize it to be a unit vector, retaining only the physically relevant direction information contained in the LRL vector. This will be called the LRL direction vector.
@@
The rest of the paper is organized as follows. In section 2, the generators of the infinitesimal symmetry transformations given by Noether’s theorem in reverse applied to the constants of motion in the Kepler problem will be reviewed using a modern formulation. The symmetry algebras respectively containing the LRL vector and the LRL direction vector will be summarized in section 3 through the corresponding Poisson brackets by familiar use of a Hamiltonian formulation. In addition, the action of the infinitesimal symmetries on the constants of motion will be derived in terms of the Poisson brackets. In section 4, the explicit symmetry transformation group arising from the LRL direction vector will be worked out on the kinematical variables. A key part of the derivation involves the use of the infinitesimal symmetry action on the energy, angular momentum, and the LRL direction vector. In section 5, the derivation is extended to obtain the explicit form of the LRL symmetry transformation group on the kinematic variables. This turns out to be more complicated because the structure of the Poisson brackets involving the LRL vector is richer in comparison to the Poisson brackets which involve the LRL direction vector. Some concluding remarks will be given in section 6.
@@ 2. Infinitesimal symmetry generators
where $\kappa$ is a positive constant. For particle motion in Newtonian gravity, $\vec{r}$} is the position of the particle with respect to the central mass, while $\kappa=GM$ where $M$ is the central mass and $G$ is the gravitational constant. For two body motion under Newtonian gravitational attraction, $\vec{r}$} is the relative position of the bodies with respect to their center of mass, and $\kappa=G\mu$ where $\mu$ is the reduced mass of the two bodies. In both cases, $\hat{r}=|{\vec{r}}|^{-1}{\vec{r}}$ is the unit position vector.
@@
The LRL vector is orthogonal to the angular momentum, $\vec{A}\cdot\vec{L}=0$, and has magnitude $|\vec{A}|=\sqrt{\kappa^{2}+2E|\vec{L}|^{2}}$. For circular orbits, $\vec{A}$ vanishes, whereas for elliptical, parabolic, and hyperbolic orbits, $\vec{A}$ is non-trivial and points in the direction of the periapsis. Its magnitude is proportional to the eccentricity of the orbit, $e=|\vec{A}|/\kappa$. Therefore, among the three components of $\vec{A}$, only the angle of $\vec{A}$ in the plane of motion for non-circular orbits represents an independent constant of motion. If this angle, $\Theta$, is measured relative to an axis defined by the periapsis point in the orbit, then it is given by $\Theta=0$.
Thus, $E$, $\vec{L}$, $\Theta$ provide 5 independent constants of motion. This is the maximal number for any type of motion in three dimensions. There is an additional conserved quantity, which involves $t$ and is not constant of motion, that can be defined as the time $T$ at which the periapsis point is reached in a non-circular orbit. An explicit formula for $T$, along with discussion of its physical significance, is given in Ref.[11] for general central force dynamics.
Noether’s theorem gives a one-to-one correspondence between conserved quantities and infinitesimal dynamical symmetries of an action principle. For the Kepler problem (2.1), the action principle is $S=\int_{t_{1}}^{t_{2}}\mathcal{L}({\vec{r}},\dot{{\vec{r}}})\,dt$ where
$\mathcal{L}({\vec{r}},\dot{{\vec{r}}})=\tfrac{1}{2}|\dot{{\vec{r}}}|^{2}+\kappa|{\vec{r}}|^{-1}$ (2.5)
is the Lagrangian, whose variational derivative yields the equations of motion,
$\frac{\partial{\mathcal{L}}}{\partial{{\vec{r}}}}-\frac{d}{dt}\frac{\partial{\mathcal{L}}}{\partial{\dot{{\vec{r}}}}}=-\ddot{{\vec{r}}}-\kappa|{\vec{r}}|^{-2}{\hat{r}}$. (2.6)
Recall that adding a total time derivative to $\mathcal{L}$, which just changes $S$ by the addition of an endpoint term, does not affect the equations of motion. In this context, an infinitesimal dynamical symmetry is a generator
@@
acting on the dynamical variable $\vec{r}$}, such that the action principle is invariant up to an endpoint term. Specifically, the condition of invariance is given by ${\rm pr}{\bf X}(S)=W|_{t_{1}}^{t_{2}}$ for some function $W(t,{\vec{r}},\dot{{\vec{r}}})$, where
${\rm pr}{\bf X}={\vec{P}}\cdot\partial_{{\vec{r}}}+\dot{{\vec{P}}}\cdot\partial_{\dot{{\vec{r}}}}$ (2.8)
denotes the prolongation of $\bf X$}, acting on the Lagrangian variables $\vec{r}$} and $\dot{{\vec{r}}}$. Infinitesimal symmetry invariance can be usefully restated entirely in terms of the Lagrangian as
@@
A function $C(t,{\vec{r}},\dot{{\vec{r}}})$ is a conserved quantity for the Kepler problem, namely $\dot{C}=0$ for all solutions of the equations of motion (2.1), if and only if the generator ${\bf X}={\vec{P}}(t,{\vec{r}},\dot{{\vec{r}}})\cdot\partial_{{\vec{r}}}$ is an infinitesimal dynamical symmetry of the Lagrangian (2.5), where
${\vec{P}}=\dfrac{\partial C}{\partial\dot{{\vec{r}}}}.$ (2.10)
Any infinitesimal dynamical symmetry gives rise to a (one-parameter) transformation group of dynamical symmetries which simply represents the flow generated by the vector field ${\rm pr}{\bf X}$. This can be understood in two different ways.
@@
with $\varepsilon$ being the group parameter, where $\varepsilon=0$ yields the identity transformation. Note that this transformation group keeps $t$ fixed.
A second way to define an equivalent transformation group is by having the transformations act directly on solutions $({\vec{r}}(t),{\vec{v}}(t))$ of the equations of motion (2.1). This yields
@@
where, now,
${\rm pr}{\bf X}={\vec{P}}(t,{\vec{r}}(t),{\vec{v}}(t))\cdot\partial_{{\vec{r}}}+{\vec{P}}(t,{\vec{r}}(t),{\vec{v}}(t))\dot{\mathstrut}\cdot\partial_{{\vec{v}}}$ (2.14)
using the ordinary time derivative restricted to solutions.
Both transformation groups (2.12) and (2.13) represent the one-parameter dynamical symmetry group generated by $\bf X$}.
@@
where $\vec{\xi}$ and $\tau$ are functions only of $t$ and $\vec{r}$}. Their infinitesimal generator is given by
${\bf Y}_{\text{point}}=\tau\partial_{t}+\vec{\xi}\cdot\partial_{{\vec{r}}}+\dot{\vec{\xi}}\cdot\partial_{{\vec{v}}},$ (2.16)
corresponding to the transformation group $(t,{\vec{r}},{\vec{v}})^{*}=\exp(\varepsilon\,{\bf Y}_{\text{point}})(t,{\vec{r}},{\vec{v}})$. These symmetries act in the coordinate space $(t,{\vec{r}},{\vec{v}})$, unlike dynamical symmetries which act in phase space $({\vec{r}},{\vec{v}})$ or on solutions of the equations of motion.
Nevertheless, point symmetries naturally give rise to a transformation group acting in phase space. The transformations are readily derived by considering the effect of a point symmetry on solutions of the equations of motion (2.1), where each solution geometrically describes a curve $({\vec{r}}(t),{\vec{v}}(t))$ in the coordinate space $(t,{\vec{r}},{\vec{v}})$. Under a point symmetry, the curve is mapped to another curve $({\vec{r}}^{*}(t^{*}),{\vec{v}}^{*}(t^{*}))$ in the same space. This mapping has an equivalent formulation in which $t$ is kept fixed and only $({\vec{r}}(t),{\vec{v}}(t))$ is changed: the resulting infinitesimal transformation has the form [12]
${\vec{r}}{}^{*}(t)={\vec{r}}(t)+\varepsilon\big(\vec{\xi}(t,{\vec{r}}(t))-\tau(t,{\vec{r}}(t))\,{\vec{v}}(t)\big)+O(\varepsilon^{2}),$ (2.17)
${\vec{v}}{}^{*}(t)={\vec{v}}(t)+\varepsilon\big(\vec{\xi}(t,{\vec{r}}(t))-\tau(t,{\vec{r}}(t))\,{\vec{v}}(t)\big)\dot{\mathstrut}+O(\varepsilon^{2}).$
Its corresponding infinitesimal symmetry generator has the following simple formulation in phase space,
$\hat{\bf X}_{\text{point}}={\vec{P}}_{\text{point}}\cdot\partial_{{\vec{r}}}+\mathcal{D}_{t}{\vec{P}}_{\text{point}}\cdot\partial_{{\vec{v}}},\quad{\vec{P}}_{\text{point}}=\vec{\xi}-\tau{\vec{v}}.$ (2.18)
Observe that this generator (2.18) is related to the previous generator (2.16) by ${\bf Y}_{\text{point}}=\hat{\bf X}_{\text{point}}+\tau D_{t}$, where
@@
Consequently, what distinguishes a point symmetry generator from a dynamical symmetry generator is their dependence on $\vec{v}$}.
Proposition 2.
A point symmetry generator ${\vec{P}}_{\text{point}}(t,{\vec{r}},{\vec{v}})$ is strictly linear in $\vec{v}$}, namely,
$\dfrac{\partial{\vec{P}}_{\text{point}}}{\partial{\vec{v}}}=\tau(t,{\vec{r}})\mathbf{I}$ (2.20)
where $\mathbf{I}$ denotes the identity matrix. A genuine dynamical symmetry generator ${\vec{P}}_{\text{dyn.}}(t,{\vec{r}},{\vec{v}})$ is either non-strictly linear or nonlinear in $\vec{v}$}, namely,
$\dfrac{\partial{\vec{P}}_{\text{dyn.}}}{\partial{\vec{v}}}=\mathbf{f}(t,{\vec{r}},{\vec{v}})$ (2.21)
with the matrix function $\mathbf{f}(t,{\vec{r}},{\vec{v}})$ satisfying
$\frac{\partial{\mathbf{f}}}{\partial{{\vec{v}}}}=0,\quad\mathbf{f}\neq\tau\mathbf{I}$ (2.22a)
or
$\frac{\partial{\mathbf{f}}}{\partial{{\vec{v}}}}\neq 0.$ (2.22b)
An important observation is that any infinitesimal dynamical symmetry (2.7) can be extended to act in the coordinate space $(t,{\vec{r}},{\vec{v}})$ by means of the generator
@@
with parameter $\varepsilon$. In particular,
$t{}^{*}=t+\varepsilon\tau+O(\varepsilon^{2}),\quad{\vec{r}}{}^{*}={\vec{r}}+\varepsilon({\vec{P}}+\tau{\vec{v}})+O(\varepsilon^{2}),\quad{\vec{v}}{}^{*}={\vec{v}}+\varepsilon(\mathcal{D}_{t}{\vec{P}}-\tau|{\vec{r}}|^{-2}{\hat{r}})+O(\varepsilon^{2}).$ (2.25)
In this transformation group, $\tau$ constitutes a gauge freedom, since both ${\rm pr}{\bf X}$ and $\bf Y$} act in an identical way on solutions $({\vec{r}}(t),{\vec{v}}(t))$.
Existence of this gauge freedom has not been widely recognized in the literature. However, it will turn out to play a key role in finding the explicit form of dynamical symmetry transformations generated from the LRL vector

### v3

@@ 1. Introduction
The Kepler problem in Newtonian mechanics has been widely studied and encompasses both a particle moving under inverse-square force and two bodies moving under mutual gravitational attraction. One of its most interesting features is the existence of the conserved Laplace–Runge–Lenz (LRL) vector (see e.g. [1, 2] and references therein). It lies in the plane of motion, which is orthogonal to the angular momentum vector, and points in the direction given by the periapsis in the case of inverse-square motion and the position of closest separation in the case of two-body motion.
@@
In contrast to the conservation of energy and angular momentum in the Kepler problem, which come directly from Noether’s theorem using the manifest time-translation and rotation symmetries of Newtonian space, the LRL vector does not arise from any kinematic symmetry principle. Nevertheless, Noether’s theorem in reverse (e.g. [3, 4]) states that every conserved quantity is generated by some infinitesimal symmetry transformation that leaves invariant the action principle. Since the LRL vector comprises three independent conserved quantities, its symmetry generator consists of three independent infinitesimal transformations, which were first worked out in Ref.[5]. They turn out to involve the velocity variable, unlike the infinitesimal time-translation and rotation transformations which involve only the time and position variables. In particular, the LRL infinitesimal symmetry transformations are primary examples of what are called dynamical symmetries, in contradistinction to point symmetries which are typified by infinitesimal time-translations and rotations.
@@
When the LRL symmetries are normalized to have the same physical dimension as the rotations, these 6 point symmetries are known to form a 6-dimensional Lie algebra whose structure turns out to be $\mathfrak{so}(4)$ when the energy is negative, and $\mathfrak{so}(3,1)$ when the energy is positive [6]. These Lie algebras are isomorphic to the algebra of infinitesimal isometries of, respectively, a 3-sphere $S^{3}$ in $\mathbb{R}^{4}$ and a 3-hyperboloid $H^{3}$ in Minkowski space $\mathbb{R}^{3+1}$.
@@
One explanation for the appearance of four-dimensional spaces is provided by Moser’s observation [7] that, in the negative energy case, the Kepler problem is equivalent to non-affine geodesic motion on $S^{3}$ under stereographic projection. A kinematical description of the projection is given by [8, 6] identifying the stereographic coordinates in an auxiliary $\mathbb{R}^{4}$ as the components of the Kepler velocity variable. The rotations in the auxiliary space constitute a $SO(4)$ transformation group that combines the physical $SO(3)$ rotations and the LRL symmetry transformations. A similar equivalence [9] holds in the positive energy case by considering non-affine geodesic motion on $H^{3}$.
@@
Another remaining question, which is not commonly discussed in the literature, arises from the magnitude of LRL vector which is given by a scalar expression in terms of the energy and magnitude of the angular momentum. On the one hand, this choice of normalization of the LRL vector is essential for the algebra generated by the LRL symmetries and rotations to be $\mathfrak{so}(4)$ when energy is negative and $\mathfrak{so}(3,1)$ when energy is positive. Yet on the other hand, it is a free choice, which raises the question of whether a simpler structure is possible for the symmetry algebra if some other choice of normalization is made. A natural choice is to normalize it to be a unit vector, retaining only the physically relevant direction information contained in the LRL vector. This will be called the LRL direction vector.
@@
The rest of the paper is organized as follows. In section 2, the generators of the infinitesimal symmetry transformations will be reviewed using a modern formulation. The symmetry algebras respectively containing the LRL vector and the LRL direction vector will be summarized in section 3 in terms of Poisson brackets. In addition, the action of the infinitesimal symmetries on the constants of motion will be derived. In section 4, the explicit symmetry transformation group arising from the LRL direction vector will be worked out on the kinematical variables. In section 5, the derivation is extended to obtain the explicit form of the LRL symmetry transformation group on the kinematic variables. Some concluding remarks will be given in section 6.
@@ 2. Infinitesimal symmetry generators
where $\kappa$ is a positive constant and $\hat{r}=|{\vec{r}}|^{-1}{\vec{r}}$ is a unit vector. For particle motion in Newtonian gravity, $\vec{r}$} is the position of the particle with respect to the central mass, while $\kappa=GM$ where $M$ is the central mass and $G$ is the gravitational constant. For two body motion under Newtonian gravitational attraction, $\vec{r}$} is the relative position of the bodies with respect to their center of mass, and $\kappa=G\mu$ where $\mu$ is the reduced mass of the two bodies.
@@
The LRL vector is orthogonal to the angular momentum, $\vec{A}\cdot\vec{L}=0$, and has magnitude $|\vec{A}|=\sqrt{\kappa^{2}+2E|\vec{L}|^{2}}$. For circular orbits, $\vec{A}$ vanishes, whereas for elliptical, parabolic, and hyperbolic orbits, $\vec{A}$ is non-trivial and points in the direction of the periapsis. Its magnitude is proportional to the eccentricity of the orbit, $e=|\vec{A}|/\kappa$. Therefore, among the three components of $\vec{A}$, only the angle of $\vec{A}$ with respect to a fixed axis in the plane of motion for non-circular orbits represents an independent constant of motion.
Thus, $E$, $\vec{L}$, $\Theta$ provide 5 independent constants of motion. This is the maximal number for any type of motion in three dimensions. There is an additional conserved quantity, which involves $t$ and thus is an integral of motion rather than a constant of motion, that can be defined as the time $T$ at which the periapsis point is reached in a non-circular orbit. An explicit formula for $T$, along with discussion of its physical significance, is given in Ref.[11].
Noether’s theorem gives a one-to-one correspondence between conserved quantities and infinitesimal dynamical symmetries of an action principle. For the Kepler problem (2.1), the action principle is $S=\int_{t_{1}}^{t_{2}}\mathcal{L}({\vec{r}},\dot{{\vec{r}}})\,dt$ where $\mathcal{L}({\vec{r}},\dot{{\vec{r}}})=\tfrac{1}{2}|\dot{{\vec{r}}}|^{2}+\kappa|{\vec{r}}|^{-1}$ is the Lagrangian, whose variational derivative yields the equations of motion, $\delta\mathcal{L}/\delta{\vec{r}}=-\ddot{{\vec{r}}}-\kappa|{\vec{r}}|^{-2}{\hat{r}}$. Recall that adding a total time derivative to $\mathcal{L}$, which just changes $S$ by the addition of an endpoint term, does not affect the equations of motion. In the Lagrangian setting, an infinitesimal dynamical symmetry is a generator acting on the dynamical variable $\vec{r}$},
@@
such that the action principle is invariant up to an endpoint term. Specifically, the condition of invariance is given by ${\rm pr}{\bf X}(S)=W|_{t_{1}}^{t_{2}}$ for some function $W(t,{\vec{r}},\dot{{\vec{r}}})$, where ${\rm pr}{\bf X}={\vec{P}}\cdot\partial_{{\vec{r}}}+\dot{{\vec{P}}}\cdot\partial_{\dot{{\vec{r}}}}$ denotes the prolongation of $\bf X$}, acting on the Lagrangian variables $\vec{r}$} and $\dot{{\vec{r}}}$. Infinitesimal symmetry invariance can be usefully restated entirely in terms of the Lagrangian as
@@
A function $C(t,{\vec{r}},\dot{{\vec{r}}})$ is a conserved quantity for the Kepler problem, namely $\dot{C}=0$ for all solutions of the equations of motion (2.1), if and only if the generator ${\bf X}={\vec{P}}(t,{\vec{r}},\dot{{\vec{r}}})\cdot\partial_{{\vec{r}}}$ is an infinitesimal dynamical symmetry of the Lagrangian $\mathcal{L}$, where ${\vec{P}}=\dfrac{\partial C}{\partial\dot{{\vec{r}}}}$. This states the Noether correspondence in an explicit form.
Any infinitesimal dynamical symmetry gives rise to a (one-parameter) transformation group of dynamical symmetries which geometrically represents the flow generated by the vector field ${\rm pr}{\bf X}$. This can be understood in two different ways.
@@
with $\varepsilon$ being the group parameter, where $\varepsilon=0$ yields the identity transformation.
A second way to define an equivalent transformation group is by having the transformations act directly on solutions $({\vec{r}}(t),{\vec{v}}(t))$ of the equations of motion (2.1):
@@
Both transformation groups (2.8) and (2.9) represent the one-parameter dynamical symmetry group generated by $\bf X$}. Note that $t$ is kept fixed in these transformations.
@@
where $\vec{\xi}$ and $\tau$ are functions only of $t$ and $\vec{r}$}. Their infinitesimal generator is given by ${\bf Y}_{\text{point}}=\tau\partial_{t}+\vec{\xi}\cdot\partial_{{\vec{r}}}+\dot{\vec{\xi}}\cdot\partial_{{\vec{v}}}$, corresponding to the transformation group $(t,{\vec{r}},{\vec{v}})^{*}=\exp(\varepsilon\,{\bf Y}_{\text{point}})(t,{\vec{r}},{\vec{v}})$. These symmetries act in the coordinate space $(t,{\vec{r}},{\vec{v}})$, unlike dynamical symmetries which act in phase space $({\vec{r}},{\vec{v}})$ or on solutions of the equations of motion. Nevertheless, point symmetries have an equivalent formulation in terms of a generator (2.5) acting in phase space, which arises from how they geometrically map solution curves $({\vec{r}}(t),{\vec{v}}(t))$ into themselves in the coordinate space $(t,{\vec{r}},{\vec{v}})$. This mapping, in which $t$ is kept fixed, is given by [4]
${\vec{r}}{}^{*}(t)={\vec{r}}(t)+\varepsilon\big(\vec{\xi}(t,{\vec{r}}(t))-\tau(t,{\vec{r}}(t))\,{\vec{v}}(t)\big)+O(\varepsilon^{2})$ (2.11)
along with ${\vec{v}}{}^{*}(t)=\frac{d}{dt}{\vec{r}}{}^{*}(t)$. The resulting infinitesimal symmetry generator in phase space is
$\hat{\bf X}_{\text{point}}={\vec{P}}_{\text{point}}\cdot\partial_{{\vec{r}}}+\mathcal{D}_{t}{\vec{P}}_{\text{point}}\cdot\partial_{{\vec{v}}},\quad{\vec{P}}_{\text{point}}=\vec{\xi}-\tau{\vec{v}},$ (2.12)
which is related to the previous generator ${\bf Y}_{\text{point}}=\hat{\bf X}_{\text{point}}+\tau D_{t}$, where
@@
An important corollary observation is that any infinitesimal dynamical symmetry (2.5) can be extended to act in the coordinate space $(t,{\vec{r}},{\vec{v}})$ by means of the generator
@@
with parameter $\varepsilon$. Here $\tau$ constitutes a gauge freedom, since both ${\rm pr}{\bf X}$ and $\bf Y$} act in an identical way on solutions $({\vec{r}}(t),{\vec{v}}(t))$. Existence of this gauge freedom has not been widely recognized in the literature. However, it will turn out to play a key role in finding the explicit form of dynamical symmetry transformations generated from the LRL vector. (Index notation will be adopted hereafter, with raising and lowering via the Euclidean metric $\delta_{ij}$.)